# Human-Like Robot Impedance Regulation Skill Learning from Human-Human Demonstrations

### v1

@@ Abstract
Humans are experts in collaborating with others physically by regulating compliance behaviors based on the perception of their partners’ states and the task requirements. Enabling robots to develop proficiency in human collaboration skills can facilitate more efficient human-robot collaboration (HRC). This paper introduces an innovative impedance regulation skill learning framework for achieving HRC in multiple physical collaborative tasks. The framework is designed to adjust the robot compliance to the human partner’s states while adhering to reference trajectories provided by human-human demonstrations. Specifically, electromyography (EMG) signals from human muscles are collected and analyzed to extract limb impedance, representing compliance behaviors during demonstrations. Human endpoint motions are captured and represented using a probabilistic learning method to create reference trajectories and corresponding impedance profiles. Meanwhile, an LSTM-based module is implemented to develop task-oriented impedance regulation policies by mapping the muscle synergistic contributions between two demonstrators. Finally, we propose a whole-body impedance controller for a human-like robot, coordinating joint outputs to achieve the desired impedance and reference trajectory during task execution. Experimental validation was conducted through a collaborative transportation task and two interactive Tai Chi pushing hands tasks, demonstrating superior performance from the perspective of interactive forces compared to a constant impedance control method.
Index Terms:
Impedance regulation skill learning, human-human demonstrations, human-like collaborative robot, whole-body impedance control
@@ I Introduction
Collaborative robots (cobots) have emerged as a solution for more efficient human-robot collaboration (HRC) in both industrial and domestic scenarios. Co-manipulation outperforms fully robotic manipulation by offering enhanced flexibility and effectiveness while surpasses fully human manipulation by reducing labor costs, maintaining concentration, and minimizing errors due to fatigue [1]. Consider, for instance, the various collaborative tasks in Fig. 1, distinct physical collaboration skills are required for cobots to master and apply in the form of impedance regulation skills. During human-robot co-transportation of bulky objects, humans take on the role of leaders for multiple operations while cobots undertake physical work and regulate impedance based on the human compliance behaviors, reducing physical costs and speeding up processes [2]. Tai Chi pushing hands, similar in principle to collaborative sawing tasks, requires cobots to produce inverse compliance behaviors compared to the human partner. Regardless, it remains challenging for cobots to swiftly acquire task-specific physical collaboration skills through traditional model-based methods and effectively apply them in human-robot collaboration [3].
@@
Figure 1: Illustration of human-robot physical collaborative tasks. In these scenarios, the cobot should be able to perform different compliance behaviors accordingly because physical interaction exists between the human and the cobot. We expect to learn from the human-human demonstrations and apply to a human-like mobile cobot. (We declare that the individuals in this figure are both authors and they gave permission for the use of their image.)
@@
Humans excel at collaborating with others and the kinematic and dynamic information generated during these collaborations can serve as a crucial data source. Learning from Demonstration (LfD) is an effective approach that enables cobots to acquire and imitate these skills for HRC [4]. Most LfD approaches involve a single human instructing a cobot through guided collaboration, commonly referred to as kinesthetic teaching to achieve collaborative object carrying [5] and scarf hanging [6]. While these approaches have the advantage of directly obtaining joint values during demonstrations, they prove challenging to apply to highly redundant cobots and efficiently acquire large volumes of data [7]. Other studies concentrate on learning from human-human demonstrations, which can provide extensive kinematic and dynamic information, such as reference trajectories [8] or impedance profiles [9]. Although the acquired data contains rich human behavior information, how to represent these physical behaviors and reproduce them by cobots in HRC is still an open problem [10].
@@
Trajectory and impedance are widely applied to represent human kinematic and dynamic behaviors during demonstrations. On the one hand, Gaussian Mixture Model (GMM) and Dynamic Movement Primitives (DMP) are frequently used to represent complex human demonstration motions by few parameters that can generate reference trajectories for the cobots [11]. The Task-parameterized Gaussian Mixture Model (TP-GMM) represents demonstration motions from several different perspectives for integrating task information, which allows the cobots to adjust trajectories to new situations [12]. On the other hand, stiffness and damping parameters are commonly applied to characterize the physical behavior of humans. The cobot impedance variables can be inferred based on the dynamic interaction model and learned Cartesian trajectories with optimization methods. These kinds of approaches have been successfully applied to various collaborative tasks, such as object assembly, object lifting [13], cucumbers cutting [14], and physical therapy for optimal path planning [15]. Regardless, these approaches acquire data from kinesthetic demonstration and the optimized impedance variations of the cobots may not align with those encoded physical behaviors during human-human demonstrations. Other approaches utilize bio-signals, such as electromyography (EMG), to directly record human muscle bioelectrical activity, thereby offering valuable insights into human impedance. An impedance variation approach based on human EMG feedback has been proposed to enable a cobot to achieve an appropriate impedance profile in collaborative sawing tasks [16]. Another approach has been raised for adapting the impedance of cobots based on the operator’s intentions and movements, which are estimated using a real-time EMG-to-force model [17]. However, current approaches primarily concentrate on extracting human impedance variation during collaboration while overlooking the impedance regulation based on their partners’ states, which could serve as a potential policy in human-human collaboration and be transferred to human-robot collaboration.
A variable impedance controller (VIC) is required to reproduce the learned impedance regulation skills by modulating the desired stiffness and damping of the cobot [18]. Force-based VIC methods have been proposed to achieve intuitive co-manipulation with redundant manipulators, KUKA LWR4 [19], [20]. An optimization-based method has been raised for efficient human-robot collaborative assembly with a Franka EMIKA Panda manipulator [21]. Some variable whole-body impedance control methods have been proposed for human-humanoid collaboration like ARMAR-6 [22] and Rollin’ Justin [23]. Among these methods, the human-in-the-loop VIC method is crucial for improving safety and efficiency in HRC. For instance, an online stiffness estimation based VIC method has been raised for robust collaborative polishing and assembly [24]. Another biosignal-based VIC method achieves real-time cobot stiffness regulation with the corresponding reactions of the cobot modeled with the experience/observation [16]. Nevertheless, usability and human perception should be considered or learned for designing a collaborative controller [25]. The above issues can be solved by proposing a novel VIC based on the impedance regulation skills learned from human-human demonstrations since the regulation policies can be extracted from teaching data while usability is guaranteed. Specifically, the physical reactions from the collaborator to the partner can be represented by the proper impedance regulation along the reference trajectories, which will be further transferred to the cobot for VIC development.
@@
In this article, we propose a novel skill-learning framework designed to enable cobots to acquire collaboration skills based on human-human demonstrations. Diverging from prevalent methodologies, this framework extracts the physical behaviors exhibited by human demonstrators and assimilates the proper compliance responses based on the partner’s dynamical states in collaborative tasks. Consequently, the cobot can be applied to execute a reference trajectory derived from task demonstrations and to simultaneously adjust compliance in response to the feedback of the human partner with a variable impedance controller. Two collaborative tasks including object co-transportation and confrontational Tai Chi pushing hands are raised for validation.
@@
    An impedance regulation skill learning framework is proposed for cobots to acquire collaborative compliance behaviors of humans. The TP-LQT method is applied for both motion and impedance representation from human demonstrations. Specifically, TP-GMM is adopted to represent demonstrated motion with trajectory and impedance profiles, which are extendable to new situations. Linear Quadratic Tracking (LQT) based method is then applied for cobot motion and impedance planning, which novelly corresponds compliance behaviors to the reference trajectories.
@@
    An LSTM-based impedance regulation module is constructed to innovatively map the compliance variation during human-human demonstrations. This module can dynamically adjust the cobot regulatory factor online, using EMG feedback from the partner to scale the impedance profile generated from the TP-LQT. The human compliance behaviors are represented by endpoint impedance, which is calculated with an EMG-based upper limb dynamic model.
@@
    A novel whole-body impedance controller has been designed for our specific human-like mobile cobot with a self-designed torso and two manipulators included, which can generate desired stiffness and damping during movement in Cartesian space while coordinating different components through proper torque commands.
@@ II Impedance Regulation Skills
This section presents the relationship between impedance regulation skills and physical collaboration skills with two typical modes in human-human collaboration (as shown in Fig. 2) and the definition of impedance regulation skills is given. An EMG-based stiffness estimation method is then introduced for human compliance representation, which is the preliminary of learning impedance regulation skills in demonstrations.
@@ II-A Definition
Impedance regulation skills represent a foundational aspect of physical collaboration skills at a low level. Physical collaboration skill has been defined as regulating the exchange of mechanical energy [26] or seamless adaptation of autonomy level [27] based on tasks and a mutual understanding of the partner’s intentions. To conclude, physical collaboration skills involve generating appropriate behaviors based on comprehensive feedback from both the partners and the environmental conditions. For example, during a collaborative sawing task shown in Fig. 2, one agent may demonstrate stiffly to the task by vigorously pulling the saw at one stage while exhibiting adaptability by gently pushing the saw back at another stage in response to the partner’s actions [28]. Consequently, physical collaboration skills can be conceptualized as regulatory compliance behaviors tailored to meet both task requirements and partner feedback.
@@
Figure 2: Schematic diagram of two typical physical collaboration skills and the corresponding impedance regulation skills. Leader-follower mode is widely applied in some collaborative tasks such as tug of war and collaborative transportation. The follower’s physical behaviors are concluded as performing concerted impedance according to the leader’s compliance behaviors. Mutual adaptation mode is observed in other tasks like collaborative sawing and confrontational Tai Chi, in which roles are changing during these tasks and the physical behaviors of collaborators are defined as performing adaptable impedance based on the partners.
@@
Representing human compliance behaviors can teach cobots the corresponding impedance regulation skills from human-human demonstrations since they are the most efficient capacities to transfer to the robots. Human-like variable impedance skills have been successfully learned by human-like cobots with learning frameworks through physical human-robot interfaces [29]. Nevertheless, these skills are mostly applied in the manipulation and teleoperation scenarios, in which the human compliance behaviors are directly transferred to the cobots whether offline or online [30]. In this paper, our primary focus centres on the ability to regulate behaviors in response to dynamic feedback from collaboration partners [31]. Specifically, we aim to map human impedance regulation skills that are influenced by the partner’s compliance behaviors during task execution. We select two quintessential physical collaboration skills of humans, which are targeted for replication.
Leader-follower mode: Followers are expected to align closely with leaders, contributing collaboratively and efficiently in joint tasks. Consequently, collaborators will demonstrate consistent compliance behavior relative to their partners throughout the engagement.
Mutual adaptation mode: Collaborators are expected to adapt their roles at various stages of the task, thereby exhibiting adaptable compliance in relation to their partners.
@@ II-B EMG-Based Stiffness Estimation
Human arm endpoint stiffness should be estimated to represent compliance behaviors and further acquire impedance regulation skills during demonstrations. Stochastic perturbation techniques are one of the most reliable methods to directly construct the relationship between force and displacement and therefore identify the characteristics of the human endpoint stiffness in Cartesian space. These methods are not feasible to be directly applied in human-human demonstrations or human-robot collaboration since real-time estimation is acquired. Hence, current researchers have designed a calibration process to extract the stiffness-related biosignal features (e.g., EMG signals) for estimating arm endpoint stiffness [32]. Firstly, we construct the positive-definite endpoint stiffness matrix through a geometric approach based on the process proposed in [33]:
@@
Figure 3: The overall framework scheme from human-human demonstration to human-robot collaboration. Human-Human Demonstration: multi-modal data including the 3D pose of human joints, EMG signals, and interactive forces, are recorded. By calculating the stiffness matrices through the endpoint stiffness estimation module and reference trajectories through the interaction model, the motion/impedance profiles are learned through TP-LQT. Human-Robot Collaboration: the estimated impedance profiles are sent to construct an LSTM-based impedance regulation module, which can generate proper regulatory factors during online human-robot collaboration. Finally, the reference trajectories and impedance are sent to our human-like cobot, CURI, with a whole-body impedance controller for task execution. (We declare that the individuals in this figure are both authors and they gave permission for the use of their image.)
@@
We define that vector $\bm{l}_{1}$ starts from the shoulder joint to the elbow joint and vector $\bm{l}_{2}$ starts from the shoulder joint to the wrist joint. $\bm{l}_{3}$ is a perpendicular vector starts from the elbow joint to the vector $\bm{l}_{1}$ while $\bm{l}_{4}$ represents a vector which is perpendicular to the arm triangle plane. Hence, the major, median, and minor principal axis $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ of the endpoint stiffness ellipsoid are determined as $\bm{l}_{2}$, $\bm{l}_{3}$, and $\bm{l}_{4}$. The matrices $\bm{V}\in\mathbb{R}^{3}$ and $\bm{H}\in\mathbb{R}^{3}$ can be written as:
$\bm{V}=\left[\begin{matrix}\ \displaystyle{\frac{\bm{l}_{2}}{||\bm{l}_{2}||}}\ \displaystyle{\frac{(\bm{l}_{2}\times\bm{l}_{1})\times\bm{l}_{2}}{||\bm{l}_{2}\times\bm{l}_{1})\times\bm{l}_{2}||}}\ \displaystyle{\frac{\bm{l}_{2}\times\bm{l}_{1}}{||\bm{l}_{2}\times\bm{l}_{1}||}}\ \end{matrix}\right],$ (3)
@@
where $\alpha(A)$ represents the synergistic contribution of the muscle co-contractions with $\alpha(A)=b_{1}A+b_{2}$, $A$ denotes the comprehensive muscle activation level; $\bm{H}_{s}\in\mathbb{R}^{3}$ is the diagonal matrix which contains the configuration dependence eigenvalues. Their values are given by the length of vector $l_{2}$ and $l_{3}$:
$d_{1}=||\bm{l}_{2}||,\ \ d_{2}=||\bm{l}_{3}||=||\bm{l_{1}}\cdot\frac{(\bm{l}_{2}\times\bm{l}_{1})\times\bm{l}_{2}}{||\bm{l}_{2}\times\bm{l}_{1})\times\bm{l}_{2}||}||.$ (5)
@@
where $A_{BB}$ and $A_{TB}$ represent the muscle activation levels of BB and TB, respectively. Assume that the EMG signals from the demonstrator are collected as $[U_{BB},U_{TB}]\in\mathbb{R}^{2}$ with $2$ EMG sensors in measurement at each time step. They are further processed by filters and normalized using Maximal Voluntary Contraction (MVC). Then, the muscular activation levels based on the MVC and the processed EMG data can be calculated as $[A_{BB},A_{TB}]\in\mathbb{R}^{2}$.
The parameters $a_{1}$, $a_{2}$, $b_{1}$, and $b_{2}$, which vary among individuals, are identified using an optimization method that minimizes the error between the estimated and actual endpoint stiffness matrices. A perturbation method is employed to measure the matrices [30]. Upon identifying these parameters for each demonstrator, we can estimate their endpoint stiffness matrices during demonstrations and learn representations of impedance regulation skills.

### v2

@@ Abstract
Humans are experts in physical collaboration by leveraging cognitive abilities such as perception, reasoning, and decision-making to regulate compliance behaviors based on their partners’ states and task requirements. Equipping robots with similar cognitive-inspired collaboration skills can significantly enhance the efficiency and adaptability of human-robot collaboration (HRC). This paper introduces an innovative Human-Inspired Impedance Regulation Skill Learning framework (HI-ImpRSL) for robotic systems to achieve leader-follower and mutual adaptation in multiple physical collaborative tasks. The proposed framework enables the robot to adapt its compliance based on human states and reference trajectories derived from human-human demonstrations. By integrating electromyography (EMG) signals and motion data, we extract endpoint impedance profiles and reference trajectories to construct a joint representation via imitation learning. An LSTM-based module then learns task-oriented impedance regulation policies, which are implemented through a whole-body impedance controller for online impedance adaptation. Experimental validation was conducted through collaborative transportation, two interactive Tai Chi pushing hands, and collaborative sawing tasks with multiple human subjects, demonstrating the ability of our framework to achieve human-like collaboration skills and the superior performance from the perspective of interactive forces compared to four other related methods.
@@ I Introduction
Collaborative robots (cobots) have emerged as a practical route to more intuitive and efficient human-robot collaboration (HRC) across industrial and domestic settings. By combining human adaptability with robotic precision, collaborative manipulation improves task flexibility and effectiveness while alleviating human workload and fatigue-related errors [sheng2025human]. Consider the various collaborative tasks in Fig. 1, distinct collaboration skills are required for cobots, which operate at the cognitive level and are embodied physically through impedance regulation skills. For instance, during human-robot transportation of bulky objects, humans take on the role of leaders for multiple operations while cobots undertake physical work and regulate impedance based on the human compliance behaviors, reflecting cognitive processes such as understanding leader intentions and dynamically adapting to varying compliance [nemec2016bimanual]. Tai Chi pushing hands tasks require cobots to produce inverse compliance behaviors compared to the human partner in different stages, highlighting not only the need for precise physical actions but also the cognitive ability to recognize partner strategies and adapt to changing interaction dynamics. Nevertheless, it remains challenging for cobots to swiftly acquire these impedance regulation skills through traditional model-based methods and effectively apply them in human-robot collaboration [8907351].
@@
Figure 1: Illustration of human-robot physical collaborative tasks. In these scenarios, the cobot should be able to perform different compliance behaviors accordingly because physical interaction exists between the human and the cobot. We expect to learn from the human-human demonstrations and apply them to a human-like cobot.
@@
Learning from Demonstration (LfD) is an effective approach that enables cobots to acquire and imitate these skills for HRC [ravichandar2020recent]. Most LfD approaches involve a single human instructing a cobot through guided collaboration, commonly referred to as kinesthetic teaching to achieve tasks such as collaborative object carrying [rozo2016learning] and scarf hanging [7139990]. Other studies concentrate on learning from human-human demonstrations, which directly capture physical behaviors (e.g., motion, compliance) of one partner and record relevant kinematic data [7989334] and dynamic data [zhang2022electromyography].
@@
Trajectory and impedance are widely used to represent human kinematic and dynamic behaviors [wu2024simultaneous]. On the one hand, Gaussian Mixture Model (GMM) and Dynamic Movement Primitives (DMP) are frequently used to represent complex human demonstration motions with a few parameters that can generate reference trajectories for cobots [CalinonLee19]. The Task-Parameterized Gaussian Mixture Model (TP-GMM) represents demonstration motions from several different perspectives for integrating task information, which allows the cobots to adjust trajectories to new situations [calinon2016tutorial]. On the other hand, stiffness and damping parameters are commonly applied to characterize the physical behavior of humans. The cobot impedance variables can be inferred based on the dynamic interaction model and learned Cartesian trajectories with optimization methods. Such approaches have been successfully applied to various collaborative tasks, such as object assembly [jung2024touch], object lifting [yang2018dmps], sawing [cao2023passive], and drag trajectory tracking motion in open space [zhang2025intention]. Regardless, these approaches acquire data from kinesthetic demonstration and the optimized impedance variations of the cobots may not align with those encoded physical behaviors during human-human demonstrations. Other approaches utilize biosignals, such as electromyography (EMG), to directly record human muscle bioelectrical activity, thereby offering valuable insights into human impedance [luo2024impedance]. An EMG-based adaptive variable impedance control has been validated in collaborative sawing and wiping [wang2025adaptive]. Another framework has adapted robot trajectories with force feedback, AR motion prediction, and EMG-based impedance to reduce interaction force and maintain tracking accuracy [luo2024physical]. However, these methods focus on replicating physical behaviors without considering their dependence on the partner’s state and task dynamics, which limits task generalization due to neglecting adaptive regulation. To address this, our study highlights learning impedance policies driven by cognitive principles of different collaboration skills.
A variable impedance controller (VIC) is required to reproduce the learned impedance regulation skills by modulating the desired stiffness and damping of the cobot [sharifi2021impedance]. Among these methods, the human-in-the-loop VIC is crucial for improving safety and efficiency in HRC [wang2023based]. For instance, a neuroadaptive impedance control with estimation of human-robot interaction (HRI), EMG-driven stiffness, prescribed performance, and neural compensation enables natural stable interaction with improved accuracy [liu2025neuroadaptive]. Nevertheless, usability and human perception should be considered or learned for designing a collaborative controller [maccarini2022preference]. The above issues can be solved by proposing a novel VIC based on the impedance regulation skills learned from human-human demonstrations since the regulation policies can be extracted from teaching data while usability is inherently ensured. By integrating cognitive-inspired impedance regulation with imitation learning methods, the proposed VIC framework enables robots to adapt dynamically to their human partners, contributing to the development of human-like collaboration skills learning.
@@
In this article, we propose a novel human-inspired impedance regulation skill learning framework (HI-ImpRSL) designed to enable cobots to acquire collaboration skills based on human-human demonstrations. Diverging from prevalent methodologies, HI-ImpRSL extracts the physical behaviors exhibited by human demonstrators and assimilates the proper compliance responses based on the partner’s dynamical states in collaborative tasks. Consequently, the cobot can be applied to execute a reference trajectory derived from task demonstrations and to simultaneously adjust compliance in response to the feedback of the human partner with a variable impedance controller. Three collaborative tasks, including collaborative transportation, confrontational unimanual and bimanual Tai Chi pushing hands, and collaborative sawing, are raised for validation with various subjects.
@@
    HI-ImpRSL framework is proposed for cobots to acquire collaborative compliance behaviors of humans. The imitation learning method is applied for joint motion-impedance representation from human demonstrations. Specifically, TP-GMM is adopted to represent demonstrated trajectory and impedance profiles, which are extendable to new situations (e.g., start-goal configurations, collaborator kinematics). The Linear Quadratic Tracking (LQT) based method is then applied for cobot motion and impedance planning, which explicitly couples compliance behaviors with the reference trajectories.
@@
    An LSTM-based impedance regulation module is constructed to map the compliance variation during human-human demonstrations. This module can dynamically adjust the cobot regulatory factor online, using EMG feedback from the partner to scale the learned impedance profile. The generalization of this module to different subjects has been verified through both human-human and human-robot collaboration.
@@
    A novel whole-body impedance controller has been designed for our human-like mobile cobot with a self-designed torso and two manipulators included, which can generate desired stiffness and damping during movement in Cartesian space while coordinating different components through proper torque commands.
  • 4)
    A comprehensive comparison with four other methods shows the superior performance of the HI-ImpRSL by collaborating with various human subjects.
@@ II Impedance Regulation Skills
This section presents the relationship between impedance regulation skills and collaboration skills with two typical modes in human-human collaboration (as shown in Fig. 2), and the definition of impedance regulation skills is given. An EMG-based stiffness estimation method is then introduced for human compliance representation, which is the preliminary of learning impedance regulation skills in demonstrations.
@@ II-A Definition
Impedance regulation skills represent a foundational aspect of collaboration skills at a low level. Collaboration skill has been defined as regulating the exchange of mechanical energy [5379513] or seamless adaptation of autonomy level [9501975] based on tasks and a mutual understanding of the partner’s intentions. To conclude, collaboration skills involve generating appropriate behaviors based on comprehensive feedback with cognitive processes. For example, during a collaborative sawing task shown in Fig. 2, one agent may demonstrate stiffly in the task by vigorously pulling the saw at one stage while exhibiting adaptability by gently pushing the saw back at another stage in response to the partner’s actions [ZENG2021103668]. Consequently, collaboration skills can be conceptualized as regulatory compliance behaviors tailored to meet both task requirements and partner feedback at physical levels.
@@
Figure 2: Schematic diagram of two typical collaboration skills and the corresponding impedance regulation skills. Leader-follower: the follower’s physical behaviors are concluded as performing concerted impedance according to the leader. Mutual adaptation: roles are changing during these tasks and the physical behaviors of collaborators are defined as performing adaptable impedance based on the partners.
@@
Representing human compliance behaviors can teach cobots the corresponding impedance regulation skills from human-human demonstrations since they are the most efficient capacities to transfer to the robots. Human-like impedance regulation skills have been successfully learned by human-like cobots with learning frameworks through physical human-robot interfaces [6224904]. Nevertheless, these skills are mostly applied in the manipulation and teleoperation scenarios, in which the human compliance behaviors are directly transferred to the cobots whether offline or online [arashijrr]. In human-human collaboration, two quintessential physical collaboration modes, leader-follower and mutual adaptation, are widely observed, and their differences serve as a comprehensive foundation for studying the regulation of compliance behaviors (Fig. 2).
Leader-follower mode: Followers are expected to align closely with leaders from the perspective of compliance behaviors. Consequently, collaborators will demonstrate analogous endpoint impedance variations relative to their partners throughout the task execution. For example, during a collaborative transportation task, one individual leads the movement, and the other adapts their compliance to follow and support the shared goal. This mode is characterized by unilateral compliance regulation, where the follower reacts to the leader’s compliance to ensure smooth and efficient collaboration.
Mutual adaptation mode: Collaborators are expected to adapt their roles at different stages of the task, thereby exhibiting adaptable compliance in relation to their partners. For instance, in Tai Chi pushing hands, both participants continuously adjust their compliance and interaction forces in response to each other’s movements. This mode requires bidirectional compliance regulation and dynamic role switching, which emphasizes mutual understanding and adaptation.
These two modes highlight distinct collaboration dynamics: leader-follower tasks focus on reactive compliance, while mutual adaptation tasks require simultaneous proactive and reactive behaviors. Studying these modes provides insight into how humans regulate their compliance based on both task demands and their partner’s actions, offering a foundation for understanding physical collaboration.
@@ II-B EMG-Based Stiffness Estimation
Human arm endpoint stiffness should be estimated to represent compliance behaviors and further acquire impedance regulation skills during demonstrations. Stochastic perturbation techniques are one of the most reliable methods to directly construct the relationship between force and displacement and therefore identify the characteristics of the human endpoint stiffness in Cartesian space. These methods are not feasible to be directly applied in human-human demonstrations or human-robot collaboration since real-time estimation is acquired. Hence, current researchers have designed a calibration process to extract the stiffness-related biosignal features (e.g., EMG signals) for estimating arm endpoint stiffness [osu1999]. Firstly, we construct the positive-definite endpoint stiffness matrix through a geometric approach based on the process proposed in [s20185357]:
@@
We define that vector $\bm{l}_{1}$ starts from the shoulder joint to the elbow joint and vector $\bm{l}_{2}$ starts from the shoulder joint to the wrist joint. $\bm{l}_{3}$ is a perpendicular vector starts from the elbow joint to the vector $\bm{l}_{1}$ while $\bm{l}_{4}$ represents a vector which is perpendicular to the arm triangle plane. Hence, the major, median, and minor principal axis $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ of the endpoint stiffness ellipsoid are determined as $\bm{l}_{2}$, $\bm{l}_{3}$, and $\bm{l}_{4}$. The matrices $\bm{V}\in\mathbb{R}^{3\times 3}$ and $\bm{H}\in\mathbb{R}^{3\times 3}$ can be written as:
[…]
@@
where $\alpha(A)$ represents the synergistic contribution of the muscle co-contractions with $\alpha(A)=b_{1}A+b_{2}$, $A$ denotes the comprehensive muscle activation level, $b_{1}$ and $b_{2}$ are the model parameters to be identified; $\bm{H}_{s}\in\mathbb{R}^{3\times 3}$ is the diagonal matrix that also contains the personalized parameters $a_{1},a_{2}$ and configuration dependence eigenvalues. Their values are given by the length of vector $l_{2}$ and $l_{3}$:
[…]
Figure 4: Multi-modal data acquisition records synchronized human motion and interaction signals, taking as inputs the joint/endpoint poses, EMG-based muscle activations, and interaction forces. Endpoint stiffness estimation uses human kinematics and muscle activation levels as inputs to estimate the human endpoint stiffness, providing impedance profiles for learning and regulation. The interaction model takes the measured TCP trajectory, the estimated stiffness, and the interaction force as inputs to compute the human reference trajectory consistent with a spring–damper relationship. The TP-LQT module consumes the human reference trajectory, the human stiffness profile, and task parameters to generate the robot reference trajectory and the robot stiffness profile. The LSTM-based impedance regulation module takes the collaborator’s muscle co-contraction as input and outputs a robot impedance scaling index for online adaptation. Finally, the whole-body impedance controller receives the robot reference trajectory, the current robot TCP trajectory, and the reference impedance parameters as inputs and outputs joint torques for the arm and torso to execute the task on CURI.
@@
where $A_{BB}$ and $A_{TB}$ represent the muscle activation levels of BB and TB, respectively. Assume that the EMG signals from the demonstrator are collected as $U_{BB},U_{TB}]\in\mathbb{R}^{2}$ with $2$ EMG sensors in measurement at each time step. They are further processed by filters and normalized using Maximal Voluntary Contraction (MVC). Then, the muscular activation levels based on the MVC and the processed EMG data can be calculated as $A_{BB},A_{TB}]\in\mathbb{R}^{2}$.
The parameters $a_{1}$, $a_{2}$, $b_{1}$, and $b_{2}$, which vary among individuals, are identified using an optimization method that minimizes the error between the estimated and actual endpoint stiffness matrices. A perturbation method is employed to measure the matrices [arashijrr]. Upon identifying these parameters for each demonstrator, we can estimate their endpoint stiffness matrices during demonstrations and learn representations of impedance regulation skills.
II-C Interaction Force-Based Stiffness Estimation
In addition to EMG-based stiffness estimation, the end-effector impedance can be observed online based on the interaction force and position. Since offline identification is sensitive to stochastic disturbances, the stiffness-damping model can be regarded as a latent state and updated via an Extended Kalman Filter (EKF), which linearises the interaction dynamics and fuses force–position residuals for real-time parameter adaptation. By referring to the observer design in [zhang2023observer], a state vector comprising Cartesian stiffness and damping, along with a measurement model, can achieve the stiffness estimation.